OpenStudy (ibeyoshiii):

May someone graph -4x + 3y = 6?

OpenStudy (anonymous):

well first you have to get it into y=mx+b form do you know how to do that?

OpenStudy (ibeyoshiii):

Unfortunately, no. :/

OpenStudy (anonymous):

ok well here is the formulas answer y=4x 3 +2 so now you want to graph it right?

OpenStudy (anonymous):

ops it is 4x/3+2 sorry

OpenStudy (ibeyoshiii):

Yup.

OpenStudy (anonymous):

ok so do first of are you familiar with rise/run and what y=mx+b means?

OpenStudy (ibeyoshiii):

I know rise and run but the y=mx+b is what I don't really know very well.

OpenStudy (ibeyoshiii):

rise over run*

OpenStudy (anonymous):

ok well the m=the slope x=the variable b=the y-intercept do you get that so far? :)

OpenStudy (ibeyoshiii):

Yes, I do. :D

OpenStudy (anonymous):

 sorry my gragh is bad but do you get it a little better now?